OpenStudy (anonymous):

What is the value of this expression? (8 + 3)2 + 4 · (–3) 109 103 –63 –375

OpenStudy (isaiah.feynman):

Use your BEDMAS properly here.

OpenStudy (anonymous):

you mean PEMDAS

OpenStudy (isaiah.feynman):

What's P? in PEDMAS? @tatumrocks

OpenStudy (anonymous):

( )'S

OpenStudy (isaiah.feynman):

WOW high school math has changed. Lol

OpenStudy (ybarrap):

$$ (8 + 3)2 + 4 · (–3) $$ Step 1 $$ 8+3 = ? $$ Step 2 Multiply the result in step 1 by 2 Step 3 $$ 4 \times -3=? $$ Step 4 Add the result of Step 3 to that obtained from Step 2.

OpenStudy (anonymous):

PEMDAS - Parantheses exponentials multiplaction division addition subtraction

OpenStudy (anonymous):

11 and -12

OpenStudy (ybarrap):

This is the result from Step 1 and Step 3. You still need Step 2.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@DSS

OpenStudy (ybarrap):

You have the result of Step 1 as 11, right? So now for Step 2 - Multiply this result by 2 $$ (8 + 3)\bf{\color{red}{2}} + 4 · (–3) =\\ (\bf{11})\bf{\color{red}{2}} + 4 · (–3) =\ $$

OpenStudy (anonymous):

1)10 2)10

OpenStudy (ajanijones):

The answer should be ten right

OpenStudy (anonymous):

109 103 –63 –375

OpenStudy (ybarrap):

$$ (8 + 3)2 + 4 · (–3) =\\ \text{Step 1}\\ (11)2 + 4 · (–3) \\ \text{Step 2}\\ 22 + 4 · (–3) \\ \text{Step 3}\\ 22 + (-12) \\ \text{Step 4}\\ 10 $$

OpenStudy (anonymous):

Which one is it 109 103 –63 –375

OpenStudy (anonymous):

D?

OpenStudy (ybarrap):

Neither, they are all wrong. 10 is the only answer to $$ (8 + 3)2 + 4 · (–3) $$

OpenStudy (zffs):

Its actually 109. The 2 in the problem is supposed to be squared.